# Portable, Massively Parallel Implementation of a Material Point Method for Compressible Flows

### v1

@@ Abstract
The recent evolution of software and hardware technologies is leading to a renewed computational interest in Particle-In-Cell (PIC) methods such as the Material Point Method (MPM). Indeed, provided some critical aspects are properly handled, PIC methods can be cast in formulations suitable to the requirements of data locality and fine-grained parallelism of modern hardware accelerators as Graphics Processing Units (GPUs). Such a rapid and continuous technological development increases also the importance of generic and portable implementations. While continuum mechanics simulations have already shown the capabilities of MPM on a wide range of phenomena, the use of the method in compressible fluid dynamics is less frequent, especially in the supersonic regime.
In this paper we present a portable, highly parallel, GPU based MPM solver for compressible gas dynamics. The implementation aims to reach a good compromise between portability and efficiency and to give a first assessment of the potential of this approach in reproducing high speed gas flows, also taking into account solid obstacles. The proposed model constitutes a new step towards the realization of a monolithic MPM solver for Fluid-Structure Interaction (FSI) problems at all Mach numbers up to the supersonic regime.
@@
Keywords: Material Point Method; GPUs; Supersonic Gas Dynamics; Performance Portability; High-Performance Computing.
@@ 1 Introduction and Motivation
Hardware and software technology evolution has led to a renewed interest in particle-based numerical techniques for differential problems, such as the Material Point Method (MPM, [45, 56]), that was developed stemming from the Particle In Cell method Fluid Implicit Particle Method [26, 7] to treat history-dependent materials, thanks to their suitability to modern hardware accelerators requirements [25, 20]. Indeed, parallel computing is a key capability for effective advancement of industrial-grade, large-scale computations, particularly in the fields of high-speed compressible computational fluid dynamics and Fluid-Structure Interaction (FSI, [32]) simulations. On the other hand, rapid hardware and software evolution makes researchers lean towards generic and portable implementations, which enables developers to adapt fast to sudden technology innovations.
In view of exponentially increasing High Performance Computing (HPC) resources and the method’s accurate handling of a wide range of phenomena, MPM is becoming increasingly relevant for high-speed CFD and FSI simulations. In MPM, continuum media are discretized by means of Lagrangian particles, the material points, each one storing every state variable evaluated at its position, enabling the description of a wide range of materials [23, 43, 29] as well as large deformations [1, 52, 16] and fracture phenomena [53]. Notwithstanding its Lagrangian character, MPM also employs a background Cartesian grid to compute differential quantities and solve the motion equation, thus mediating the particle-particle interactions, and taking advantage of both Eulerian and Lagrangian approaches. These features make it an interesting candidate for the development of a highly parallel, Graphics Processing Units (GPUs)-based FSI solver. Indeed, in order to reach a high degree of efficiency, GPU-based solvers require a large degree of data locality and fine-level parallelism. MPM implementations can naturally achieve these, since:
@@
Since MPM first introduction, various enhancements of the method were proposed, solving some of its drawbacks such as grid-crossing error [49] and volumetric locking [15], and improving its accuracy and conservation properties. To address instabilities which arise when particles cross the grid, variants of the method were proposed, such as GIMP [4], CPDI [39], iMPM [46] and IGA-MPM [34], and numerical analysis studies on the use of B-Splines as basis functions in MPM were conducted [42]. An affine and a polynomial mapping was later added to the B-Splines MPM [30, 21] resulting in an angular momentum-conserving model. Other studies explored least squares techniques [55] enabling the recovery of the affine/polynomial splines MPM [28], the effect of spatial and temporal discretization errors, and the application of symplectic integrators [5]. In light of these developments, MPM represents an effective method for the simulation of a wide range of materials and phenomena.
@@
One of NVIDIA’s solutions for performance portability involves its own C++ compiler nvc++ [19, 37], using which one rewrites algorithm steps relying on C++ Standard Template Library (STL), specifying a parallel execution policy. To give an example, in this framework it is possible to parallelize the following for loop:
@@
Although the approach does not prevent the developer from introducing data-races or deadlocks, the code can be written only once, and then, depending on the option specified during compilation, run in parallel on multicore CPUs or GPUs.
However, nvc++ high-level features come at a price, as the developer loses the control on memory handling and, in particular, on when Host-to-Device and Device-to-Host memory copies are performed. The code runs on host, any invocation to an algorithm with a parallel execution policy expresses the developer’s preference to run that algorithm in parallel. If compiled for GPUs, memory is copied to the device, computations are performed and then data is copied back to host. In physics and in engineering dynamics simulations, where each time step depends on the previous one, this means that the maximum reachable level of parallelization corresponds to a single call to a parallel algorithm inside a traditional for loop for time advancing. Since memory fetches can be significantly more costly than computations, this is suboptimal.
The present contribution expands the scope of [2] and overcomes the performance issues with nvc++ by designing a portable MPM code for transonic and supersonic compressible gas dynamics based on Thrust [38], a C++ template library for CUDA based on the Standard Template Library (STL). In addition to nvc++ features, Thrust introduces two types of STL like vectors, one for the host (tipically, the CPU) and one for the device (tipically, the GPU); this allows the user to specify when memory should be copied via explicit calls. For example, it’s possible to perform a single Host to Device copy when reading the initial data file, and one Device To Host copy whenever saving partial results on output data files, strongly reducing execution time related to memory fetches.
@@
  • Thrust syntax is mostly compatible with STL syntax, adding just a few define preprocessor macros or templates. Thanks to this, almost the same source code can be compiled for parallel execution on GPUs with NVIDIA CUDA compiler nvcc, on AMD GPUs with the AMD compiler for ROCmThrust, and on CPUs with g++ and clang++ compilers (when it is linked with a parallel library such as TBB, Intel’s Threading Building Blocks).
@@
Notwithstanding the portability objectives, we designed the code for optimization on modern graphics hardware accelerators and made further modifications to the original algorithm [12, 58]. For example, we:
@@
Specific development choices are described that enable optimization on modern GPU-based HPC architectures and adaptability to different hardware, as well as generality of the approach for the purposes of its portability also towards different software. The accuracy and performance of the implementation are evaluated on a number of compressible flow benchmarks and results are compared with available approaches in the literature.
@@ 2.1 Model Equations and Weak Formulation
As mentioned, MPM takes advantages of both the Eulerian and the Lagrangian perspective; specifically, the Lagrangian particles discretisation approach affords flexibility in modelling various kind of phenomena and materials, while the Eulerian grid discretisation is beneficial for scalability as the grid allows the mediation of particle-particle interactions. For the mathematical modelling we choose the Lagrangian framework, and we model high speed gases as inviscid fluids in a domain $\Omega$ governed by the compressible Euler equations [24, 57]:
$\begin{cases}\dfrac{d\rho}{dt}+\rho\nabla\cdot\mathbf{v}=0\\[5.69054pt] \rho\dfrac{d\mathbf{v}}{dt}=\nabla\cdot\bm{\sigma}+\rho\mathbf{b}\\[5.69054pt] \rho\dfrac{de}{dt}=\bm{\sigma}:\bm{\dot{\varepsilon}}\end{cases}$ (1)
where $\rho$, $\mathbf{v}$, $\bm{\sigma}$, $\rho\mathbf{b}$, $p$, $e$ and $\bm{\dot{\varepsilon}}$ represent respectively the mass density, the velocity, the stress tensor, the external forces density, the pressure, the specific internal energy and the strain rate tensor of the fluid.
@@
To derive the final discretisation of the equations, a weak form of the force balance expressed by the motion equation is derived [45, 44]:
@@
where, $\mathbf{\dot{v}}$ denotes the total, or Lagrangian, velocity derivative, $\mathbf{w}$ is a generic test function and $\bm{\tau}$ the stress on the surface. It is worth noting that it is possible to derive the same formulation also by adopting the standard Finite Element procedure, provided material points are considered as moving quadrature points [17]. In fact, standard MPM can be seen as a modification of traditional FEM, in which the integration points are not related to the Finite Element mesh and are Lagrangian, and thus suitable for history dependent models too [12].
@@
In the following, we simplify further the model assuming the fluid to be inviscid, and thus the stress tensor to be due to pressure only, and we neglect external forces. Euler equations become:
The continuity and energy conservation equations will be directly discretized and used to update particle properties in the algorithm (see section 3 and algorithm 1), while the discretisation of the weak form of the motion equation is treated in more depth in the next section.
@@ 2.2 Numerical Method
The link between the Eulerian grid and the Lagrangian particles is given by the Particle-to-Grid (P2G) and Grid-to-Particle (G2P) mappings, which are the building blocks of the computational algorithm, and that can be derived from the discretization of the weak form (3). The latter can be obtained expressing the mass density via of Dirac delta functions [44]:
$\rho(\mathbf{x},t)=\sum_{k=1}^{N_{p}}M_{p}\delta(\mathbf{x}-\mathbf{x_{p}})$ (5)
Here and in the following, when it is useful to distinguish, grid quantities are written in lower case Roman letters, while particle ones are written in either upper case Roman or Greek letters.
Particle-to-Grid and Grid-to-Particle mappings read [45]:
$a=\mathrm{P2G}(A;X,Y):=\left[\sum_{k=0}^{Np-1}A_{k}\ u_{i}(X_{k},Y_{k})\right]$ (6)
$A=\mathrm{G2P}(a;X,Y):=\left[\sum_{i=0}^{(nx+1)(ny+1)-1}a_{i}\ u_{i}(X_{k},Y_{k})\right]$ (7)
$\nabla a=\mathrm{P2GD}(A;X,Y):=\left[\sum_{k=0}^{Np-1}A_{k}\ \nabla u_{i}(X_{k},Y_{k})\right]$ (8)
$\nabla A=\mathrm{G2PD}(a;X,Y):=\left[\sum_{i=0}^{(nx+1)(ny+1)-1}a_{i}\nabla u_{i}(X_{k},Y_{K})\right]$ (9)
where $u_{i}(X_{k},Y_{k})$ denotes the basis function of the solution space and test functions space centred on the $i-th$ grid node and evaluated in the $k-th$ particle position, $a$ is a generic grid variable, $A$ is its particle counterpart, $N_{p}$ is the number of particles, $nx,ny$ are the cell numbers in the $x$ and $y$ direction and $\mathrm{D}$ stands for derivative.
@@
For the numerical discretisation of the Eulerian aspects of the method, we choose a Cartesian grid of quadrilateral elements. Among the possible choices, this one lends itself naturally to computationally efficient refinements through a quad-tree approach [40]. Moreover, we opt for Lagrangian hat functions as basis for the functional spaces of both the solution and the test function, setting our treatment in a $\mathbb{Q}_{1}$ framework.
@@
This approach corresponds to bilinear weighting function of standard MPM, and is known to cause volumetric locking and spurious oscillations in the computed quantities. The reason can be easily understood by the above equations: when $u$ is (bi)-linear, $\nabla u$ is constant, but potentially different in neighbouring cells. Thus, when crossing cells, particles might experience discontinuous solicitations. This issue can be numerically tackled with many strategies, among which adding artificial viscosity [56] or choosing higher order basis functions. For simplicity of implementation, we choose the first, leaving the second for a future work.
@@
Normal unit vectors to a domain profile $\partial\Omega$ are computed as gradients of the Signed Distance Function, defined as:
@@
where $d(x,y)$ is the Euclidean distance. Taking the gradient yields:
@@
An algorithm for signed distance function computation can be found in [9].
Finally, we use explicit time integration [12], whose stability is subject to a CFL criterion:

### v2

@@ Abstract
The recent evolution of software and hardware technologies is leading to a renewed computational interest in Particle-In-Cell (PIC) methods such as the Material Point Method (MPM). Indeed, provided some critical aspects are properly handled, PIC methods can be cast in formulations suitable to the requirements of data locality and fine-grained parallelism of modern hardware accelerators as Graphics Processing Units (GPUs). Such a rapid and continuous technological development increases also the importance of generic and portable implementations. While continuum mechanics simulations have already shown the capabilities of MPM on a wide range of phenomena, the use of the method in compressible fluid dynamics is less frequent. In this paper we present a portable, highly parallel, GPU based MPM solver for compressible gas dynamics. The implementation aims to reach a good compromise between portability and efficiency and to give a first assessment of the potential of this approach in reproducing strongly compressible gas flows, also taking into account solid obstacles. The proposed model constitutes a first step towards the development of a monolithic MPM solver for Fluid-Structure Interaction (FSI) problems at all Mach numbers up to the supersonic regime.
@@
Keywords: Material Point Method; GPUs; Compressible Gas Dynamics; Performance Portability; High-Performance Computing.
@@ 1 Introduction and Motivation
Hardware and software technology evolution has led to a renewed interest in particle-based numerical techniques for differential problems, such as the Material Point Method (MPM, [51, 63]), that was developed stemming from the Particle In Cell method Fluid Implicit Particle Method [29, 8] to treat history-dependent materials, thanks to their suitability to modern hardware accelerators requirements [28, 22]. Indeed, parallel computing is a key capability for effective advancement of industrial-grade, large-scale computations, particularly in the fields of high-speed, compressible computational fluid dynamics and Fluid-Structure Interaction (FSI, [37]) simulations. On the other hand, rapid hardware and software evolution makes researchers lean towards generic and portable implementations, which enables developers to adapt fast to sudden technology innovations.
In view of exponentially increasing High Performance Computing (HPC) resources and the method’s accurate handling of a wide range of phenomena, MPM is becoming increasingly relevant for compressible CFD and FSI simulations. In MPM, continuum media are discretised by means of Lagrangian particles, the material points, each one storing every state variable evaluated at its position, enabling the description of a wide range of materials [26, 49, 32] as well as large deformations [1, 58, 18] and fracture phenomena [60]. Notwithstanding its Lagrangian character, MPM also uses a background Cartesian grid to compute differential quantities and solve the motion equation, thus mediating the particle-particle interactions, and taking advantage of both Eulerian and Lagrangian approaches. These features make it an interesting candidate for the development of a highly parallel, Graphics Processing Units (GPUs)-based FSI solver. Indeed, in order to reach a high degree of efficiency, GPU-based solvers require a large degree of data locality and fine-level parallelism. MPM implementations can naturally achieve these, since:
@@
Since MPM first introduction, various enhancements of the method were proposed, solving some of its drawbacks such as grid-crossing error [55] and volumetric locking [17], and improving its accuracy and conservation properties. To address instabilities which arise when particles cross the grid, variants of the method were proposed, such as GIMP [4], CPDI [45], iMPM [52] and IGA-MPM [40], and numerical analysis studies on the use of B-Splines as basis functions in MPM were conducted [48]. An affine and a polynomial mapping was later added to the B-Splines MPM [33, 23] resulting in an angular momentum-conserving model. Other studies explored least squares techniques [62] enabling the recovery of the affine/polynomial splines MPM [31], the effect of spatial and temporal discretisation errors, and the application of symplectic integrators [5]. In light of these developments, MPM represents an effective method for the simulation of a wide range of materials and phenomena.
@@
Therefore, existing MPM implementations either present challenges in adapting to modern hardware or are not widely studied and employed in compressible fluid dynamics. A good example is given by high-order WENO and ENO schemes, that guarantee numerically excellent results in gas dynamics. For these schemes to work properly, the stencil should be able to grow in any direction into neighbouring mesh elements. From a computational point of view, this works very well on structured grids. However, designing proper structured grids for FSI is significantly time consuming. On the other hand, if unstructured grids are chosen, growing stencils lead to a large increase of halo transfers in communication among parallel processors, thus requiring a large consumption of memory bandwidth and generating a large amount of indirection, degrading the suitability of the methods for modern parallel hardware. Moreover, such schemes have to be coupled with appropriate methods for structural mechanics, that have to be suitable for modern parallel hardware too.
In this realm, MPM presents very interesting features as complete FSI solver, indeed:
  • it benefits from the choice of employing structured Cartesian grids, since continua discretisation is delegated to Lagrangian material particles, allowing the modelling of large deformations as well;
  • it does not require coupling among solvers, as it is suitable for structural mechanics problems as well.
However, even if MPM stemmed from fluid solvers, there are only a few studies addressing compressible fluid dynamics with MPM. In addition, the existing studies, some of which (e.g. [65]) predate work on numerical linear algebra on GPUs (e.g. [6, 35]), do not deal with parallelisation on modern hardware. In addition, the MPM community has preferentially paid attention to continuum mechanics modelling.
With this work we aim to address only the fluid portion of a future FSI solver, i.e., the portion which is usually left to Eulerian CFD solvers. We focus on algorithm and data structures suitable for implementation on various parallel architectures, with a particular attention to HPC GPUs, and we address the problem of reducing limitations in treating obstacles that derives from the choice of a structured Cartesian grid, by means of a signed distance function approach.
The current study focuses on linear MPM to better study its scalability features, leaving the update to a higher order MPM for future work. However, we anticipate a discussion on how higher regularity base functions could fit our approach and how the parallel efficiency would be influenced by the extension of base function support on more elements. From the modelling point of view, to give a first assessment of MPM in simulating efficiently compressible gases on GPUs, we treat the gas as inviscid fluid, neglecting the effects of turbulence, and adopt an artificial viscosity based shock capturing technique.
@@
One of NVIDIA’s solutions for performance portability involves its own C++ compiler nvc++ [21, 43], using which one rewrites algorithm steps relying on C++ Standard Template Library (STL), specifying a parallel execution policy. To give an example, in this framework it is possible to parallelise the following for loop:
@@
Although the approach does not prevent the developer from introducing data-races or deadlocks, the code can be written only once, and then, depending on the option specified during compilation, run in parallel on multi-core CPUs or GPUs.
However, nvc++ high-level features come at a price, as the developer loses the control on memory handling and, in particular, on when Host-to-Device and Device-to-Host memory copies are performed. The code runs on host, any invocation to an algorithm with a parallel execution policy expresses the developer’s preference to run that algorithm in parallel. If compiled for GPUs, memory is copied to the device, computations are performed and then data is copied back to host. In physics and in engineering dynamics simulations, where each time step depends on the previous one, this means that the maximum reachable level of parallelisation corresponds to a single call to a parallel algorithm inside a traditional for loop for time advancing. Since memory fetches can be significantly more costly than computations, this is sub-optimal.
The present contribution expands the scope of [2] and overcomes the performance issues with nvc++ by designing a portable MPM code for transonic and supersonic compressible gas dynamics based on Thrust [44], a C++ template library for CUDA based on the Standard Template Library (STL). In addition to nvc++ features, Thrust introduces two types of STL like vectors, one for the host (typically, the CPU) and one for the device (typically, the GPU); this allows the user to specify when memory should be copied via explicit calls. For example, it’s possible to perform a single Host to Device copy when reading the initial data file, and one Device To Host copy whenever saving partial results on output data files, strongly reducing execution time related to memory fetches.
@@
  • Thrust syntax is mostly compatible with STL syntax, adding just a few define preprocessor macros or templates. Thanks to this, almost the same source code can be compiled for parallel execution on GPUs with NVIDIA CUDA compiler nvcc, on AMD GPUs with the AMD compiler for ROCm/rocThrust, and on CPUs with g++ and clang++ compilers (when it is linked with a parallel library such as TBB, Intel’s Threading Building Blocks).
@@
Notwithstanding the portability objectives, we designed the code for optimisation on modern graphics hardware accelerators and made further modifications to the original algorithm [14, 65]. For example, we:
@@
Specific development choices are described that enable optimisation on modern GPU-based HPC architectures and adaptability to different hardware, as well as generality of the approach for the purposes of its portability also towards different software. The accuracy and performance of the implementation are evaluated on a number of compressible flow benchmarks and results are compared with available approaches in the literature.
@@ 2.1 Model Equations and Weak Formulation
As mentioned, MPM takes advantages of both the Eulerian and the Lagrangian perspective; specifically, the Lagrangian particles discretisation approach affords flexibility in modelling various kind of phenomena and materials, while the Eulerian grid discretisation is beneficial for scalability as the grid allows the mediation of particle-particle interactions. For the mathematical modelling we choose the Lagrangian framework, and we model compressible fluids in a domain $\Omega$ governed by the conservation laws for mass, linear momentum and energy densities [27, 64, 19]:
$\begin{cases}\dfrac{d\rho}{dt}+\rho\nabla\cdot\mathbf{v}=0\\[5.69054pt] \rho\dfrac{d\mathbf{v}}{dt}=\nabla\cdot\bm{\sigma}+\rho\mathbf{b}\\[5.69054pt] \rho\dfrac{de}{dt}=\bm{\sigma}:\bm{\dot{\varepsilon}}-\nabla\cdot\mathbf{q}+\rho s\end{cases}$ (1)
where $\rho$, $\mathbf{v}$, $\bm{\sigma}$, $\rho\mathbf{b}$, $p$, $e$, $\bm{\dot{\varepsilon}}$, $\mathbf{q}$ and $s$ represent respectively the mass density, the velocity, the stress tensor, the external forces density, the pressure, the specific internal energy, the strain rate tensor of the fluid, the heat flux and the heat source.
@@
To obtain the final discretisation of the equations, a weak form of the force balance expressed by the motion equation is derived [51, 50]. Having chosen a proper functional space $V$ of regular enough functions, we multiply the momentum conservation equation in (1) by an arbitrary $\mathbf{w}\in V$, integrate over the fluid domain $\Omega$ and integrate by parts the divergence term.
The weak formulation of the problem then reads:
Find $\mathbf{v}$ such that, $\forall\mathbf{w}\in V$
@@
where $\mathbf{\dot{v}}$ denotes the total, or Lagrangian, velocity derivative, $\mathbf{w}$ is a generic test function and $\bm{\tau}$ is the stress on the surface. In the following, we simplify the model assuming the fluid to be inviscid, and we neglect external forces, heat sources and heat flux. We thus get the Euler equations for an inviscid, compressible fluid:
Note that, in the above expression, the fluid stress tensor is simplified from:
$\sigma_{ij}=2\mu\dot{\varepsilon}_{ij}-\tfrac{2}{3}\mu\varepsilon_{kk}\delta_{ij}-p\delta_{ij}$
to:
$\sigma_{ij}=-p\delta_{ij}$
which greatly simplifies the energy conservation equation as well as the momentum equation.
Finally, following [63], we make use of a combination of von Neumann-Richtmyer [59] and Landshoff [36] artificial viscosity as shock-capturing scheme, relying on numerical damping to handle numerical oscillations in the vicinity of discontinuities, and computing them as part of the solution to problem (4). Specifically, application of this artificial viscosity-based shock-capturing method involves the addition of an artificial term $q$ to pressure $p$, substituting the pressure $p$ to the sum $p+q$ in problem (4). The artificial term is added only in compression regions, identified by the local velocity divergence being less than zero. Its expression reads
$q=\begin{cases}c_{0}\left(\rho\,h\,\mathrm{div}(v)\right)^{2}-c_{1}\,h\,\sqrt{\gamma\tfrac{p}{\rho}}\,\mathrm{div}(v)\,\,&\mathrm{div}(v)<0\\ 0\quad&\mathrm{div}(v)\geq 0\end{cases}$ (5)
where $c_{0},c_{1}$ are dimensionless constants and $h$ is the characteristic length of a grid cell. The effects of the two components of $q$ are shown in section 4.
@@
The continuity and energy conservation equations will be directly discretised and used to update particle properties in the algorithm (see section 3 and algorithm 1), while the discretisation of the weak form of the motion equation is related to the link between the Eulerian grid and the Lagrangian particles, and is treated in more depth in the next section.
@@ 2.2 Numerical Method
The link between the Eulerian grid and the Lagrangian particles is given by the Particle-to-Grid (P2G) and Grid-to-Particle (G2P) mappings, which are the building blocks of the computational algorithm, and that can be derived during the discretisation of the weak form (3), that leads to a semi-discrete equation of the form [^1] Here and in the following, when it is useful to distinguish, grid quantities are written in lower case Roman characters, while particle ones are written in either upper case Roman or Greek characters; moreover, $i$ denotes grid nodes index, while $k$ denotes particles index.: $\forall i=0\ldots N_{h}-1\quad m_{i}\mathbf{\dot{v}}_{i}=\mathbf{f}_{i}$ (6) where $N_{h}$ denotes the number of grid nodes.
The latter can be obtained following the standard MPM procedure (see e.g. [51]), including five main steps: 1. discretise the material domain into a mesh of material sub-domains, with mass concentrated at material points $\rho(\mathbf{x},t)=\sum_{k=0}^{N_{p}-1}M_{k}\delta(\mathbf{x}-\mathbf{X}_{k})$ (7) where $N_{p}$ denotes the total number of material points, and insert (7) into (3); 2. discretise space with a standard Finite Elements Method (FEM) grid; 3. approximate any field and test function relying on FE grid shape functions $u_{i}$; 4. apply arbitrariness of test function to get the semi-discrete equation; 5. possibly, lump the mass matrix $m_{i}\!=\!\sum_{j}m_{ij}\!=\!\sum_{j}\sum_{k}M_{k}u_{i}(\mathbf{X}_{k})u_{j}(\mathbf{X}_{k})\!=\!\sum_{k}M_{k}u_{i}(\mathbf{X}_{k})$ (8) where $u_{i}(\mathbf{X}_{k})$ denotes the basis function of the solution space and test functions space centred on the $i$-th grid node and evaluated in the $k$-th particle position, $\mathbf{X}_{k}=(X_{k},Y_{k})$.
It is worth noting that it is possible to derive the same formulation also by adopting the standard Galerkin-Finite Element procedure, provided material points are considered as moving quadrature points [19]. In fact, standard MPM can be seen as a modification of traditional FEM, in which the integration points are not related to the Finite Element mesh and are Lagrangian, and thus suitable for history dependent models too [14].
Following this method, we start from the motion equation expressed in weak formulation, which reads “find $\mathbf{v}\in V$ such that equation (3) holds $\forall\mathbf{w}\in V$”, $V$ being a proper Hilbert space. To discretise the problem, we introduce a finite dimensional sub-space $V_{h}=span\{u_{i}(x,y)\}$ , project equation (3) onto $V_{h}$, and, since it has to hold for any $\mathbf{w}_{h}\in V_{h}$, we find $\mathbf{v}_{h}$ such that equation (3) holds for each function in $V_{h}$ base.
Since $V_{h}$ is finite dimensional, it is possible to express any function in terms of finite sum, e.g.
$v_{h}(x,y)=\sum_{i}v_{i}u_{i}(x,y)$ (9)
In other words, we write equation (3) choosing as test function the base functions $u_{i}$ centred on the $i-th$ grid node, and we solve numerically the integrals by means of quadrature rules, evaluating the integral at a finite set of points, called the integration points $p$, and perform the weighted sum
[…]
If the material points are taken as integration points, that is, if basis functions are evaluated in the $k-th$ particle position, with weights being their volumes (i.e. $\omega_{p}=\tfrac{M_{k}}{\rho_{k}}$), then FEM equations reduces to MPM ones. In particular,
$\mathbf{f}^{internal}_{i}=-\int_{\Omega}\bm{\sigma}\nabla u_{i}d\Omega\simeq-\sum_{k=0}^{N_{p}-1}\bm{\sigma}_{k}\tfrac{M_{k}}{\rho_{k}}\nabla u_{i}(X_{k},Y_{k})$ (11)
which is the internal component of the force in the semi-discrete equation (6), where we have already applied mass matrix lumping.
In particular, equation (11) expresses the Particle-to-Grid mapping of internal forces, which, in our inviscid gas model with $\bm{\sigma}=-p\mathbb{I}$, maps the pressure force from the material points to the grid nodes. To distinguish this from the Particle-to-Grid which does not involve moving the derivative to the test function, we call operations as (11) P2GD.
P2GD is one of the four building blocks of MPM algorithm, the others being:
  • P2G, without weak derivatives, which is needed since grid variables are reset at every time step (see sec. 3);
  • G2P, which maps quantities back to the particles after equations have been solved numerically;
  • G2PD, which also computes in-place a weak derivative.
In this work, we define these operations, which we call kernels following GPU literature, in a general way to underline their computational structure for later implementation and performance analysis, even though they are applied only as written in the algorithm in section 3. More details on how these operations are derived can be found in the original papers for the standard MPM [51], in [64, 65] for a first application to compressible fluids, and in [14, 63, 19] for later reviews.
Denote by $\mathbf{a}=[a_{i}]\in\mathbb{R}^{N_{h}}$ the array of values of some quantity $a(x,y)$ at the grid vertices $(x_{i},y_{i}),i=0\ldots N_{h}-1$, and let $\mathbf{A}=[A_{k}]\in\mathbb{R}^{N_{p}}$ be the array of values of some quantity $A(X,Y)$ at the material points $(X_{k},Y_{k}),k=0\ldots N_{p}-1$.
Particle-to-Grid and Grid-to-Particle mappings read:
$a_{i}=\mathrm{P2G}(\mathbf{A};\ x_{i},y_{i}):=\\ \sum_{k=0}^{Np-1}A_{k}\ u_{i}(X_{k},Y_{k}),\quad i=0\ldots N_{h}-1$ (12)
$A_{k}=\mathrm{G2P}(\mathbf{a};\ X_{k},Y_{k}):=\\ \sum_{i=0}^{N_{h}-1}a_{i}\ u_{i}(X_{k},Y_{k}),\quad k=0\ldots N_{p}-1$ (13)
$\left(\nabla a\right)_{i}=\mathrm{P2GD}(\mathbf{A};\ x_{i},y_{i}):=\\ \sum_{k=0}^{Np-1}A_{k}\ \left(\nabla u\right)_{i}(X_{k},Y_{k}),\quad i=0\ldots N_{h}-1$ (14)
$\left(\nabla A\right)_{k}=\mathrm{G2PD}(\mathbf{a};\ X_{k},Y_{k}):=\\ \sum_{i=0}^{N_{h}-1}a_{i}\ \left(\nabla u\right)_{i}(X_{k},Y_{K}),\quad k=0\ldots N_{p}-1.$ (15)
It is useful to generalise these kernels to the case where the mapped quantity is not stored in memory but computed on the fly as a function of other state variables, as in equation (11). To this end, let $\mathcal{F}:\mathbb{R}^{m}\to\mathbb{R}$, and rewrite the above kernels as follows :
$a_{i}=\mathrm{P2G}(\mathcal{F},\mathbf{A}^{(0)},\ldots,\mathbf{A}^{(m-1)};\ x_{i},y_{i}):=\\ \sum_{k=0}^{Np-1}\mathcal{F}\left({A}_{k}^{(0)},\ldots{A}_{k}^{(m-1)}\right)\ u_{i}(X_{k},Y_{k}),\quad i=0\ldots N_{h}-1$ (15)
$A_{k}=\mathrm{G2P}(\mathcal{F},\mathbf{a}^{(0)},\ldots,\mathbf{a}^{(m-1)};\ X_{k},Y_{k}):=\\ \sum_{i=0}^{N_{h}-1}\mathcal{F}\left({a}_{i}^{(0)},\ldots{a}_{i}^{(m-1)}\right)\ u_{i}(X_{k},Y_{k}),\;k=0\ldots N_{p}-1$ (16)
$\left(\nabla a\right)_{i}=\mathrm{P2GD}(\mathcal{F},\mathbf{A}^{(0)},\ldots,\mathbf{A}^{(m-1)};\ x_{i},y_{i}):=\\ \sum_{k=0}^{Np-1}\mathcal{F}\left({A}_{k}^{(0)},\ldots,{A}_{k}^{(m-1)}\right)\ \left(\nabla u\right)_{i}(X_{k},Y_{k}),\\ i=0\ldots N_{h}-1$ (17)
$\left(\nabla A\right)_{k}=\mathrm{G2PD}(\mathcal{F},\mathbf{a}^{(0)},\ldots,\mathbf{a}^{(m-1)};\ X_{k},Y_{k}):=\\ \sum_{i=0}^{N_{h}-1}\mathcal{F}\left({a}_{i}^{(0)},\ldots{a}_{i}^{(m-1)}\right)\ \left(\nabla u\right)_{i}(X_{k},Y_{K}),\\ k=0\ldots N_{p}-1.$ (18)
@@
For the numerical discretisation of the Eulerian aspects of the method, we choose a Cartesian grid of quadrilateral elements. Among the possible choices, this one lends itself naturally to computationally efficient refinements through a quad-tree approach [46]. Moreover, we opt for Lagrangian hat functions as basis for the functional spaces of both the solution and the test function, setting our treatment in a $\mathbb{Q}_{1}$ framework. [^2] In sec. 5 we discuss the extension to higher order basis functions.
@@
This approach corresponds to bi-linear weighting function of standard MPM, and is known to cause cell-crossing error and spurious oscillations in the computed quantities. The reason can be easily understood by the above equations: when $u$ is (bi)-linear, $\nabla u$ is constant, but potentially different in neighbouring cells. Thus, when crossing cells, particles might experience discontinuous oscillations. This issue can be numerically tackled with many strategies, among which adding artificial viscosity [63] or choosing higher order basis functions. For simplicity of implementation, we choose the first, leaving the second for a future work.
@@
Normal unit vectors to a domain profile $\partial\Omega$ are computed as gradients of the Signed Distance Function, defined as
@@
where $d(x,y)$ is the Euclidean distance. Taking the gradient yields [^3] An algorithm for signed distance function computation can be found in [10], and an example in a machine learning based framework that implicitly represents geometries from noisy surface points can be found in [13].
@@
Finally, we use explicit time integration [14], whose stability is subject to a CFL criterion